OpenStudy (anonymous):

Convert 50 mi/h to feet per minute.

OpenStudy (anonymous):

4400

OpenStudy (anonymous):

4400

OpenStudy (anonymous):

dude.

OpenStudy (accessdenied):

\[ \frac{50 \cancel{\text{ mi}}}{1\cancel{\text{ hr}}} \times \frac{5280\text{ ft}}{1 \cancel{\text{ mi}}} \times \frac{1 \cancel{\text{ hr}}}{60 \text{ min}} \] Leaves 4400 ft/min Just to explain, we're multiplying by ratios to eliminate the current unit with the new unit.

OpenStudy (anonymous):

eh, ur answer's better than mine

OpenStudy (anonymous):

google clac at its best. lmao

OpenStudy (accessdenied):

I always liked this method that my teacher would always want us to write out fully, every time we used a unit conversion. lol